OpenStudy (anonymous):

find the slope of the line that goes through the given points.(-5 9)and(-7 -5)

OpenStudy (texaschic101):

slope formula slope(m) = (y2 - y1) / (x2 - x1) (-5,9) x1 = -5, y1 = 9 (-7,-5) x2 = -7, y2 = -5 now lets sub all our stuff in m = (-5 - 9) / (-7 - (-5) m = (- 14) / (-7 + 5) m = -14 / -2 m = -14/-2 m = 7 <-- your slope any questions ?

OpenStudy (anonymous):

m = 7 <

OpenStudy (anonymous):

ok it this

OpenStudy (anonymous):

or 7

OpenStudy (johnweldon1993):

It would just be m = 7